OpenStudy (anonymous):

Solve the equation for x, accurate to three decimal places: ex − 2e−x = 1. x = 0.301 x = 0.693 x = 0.151 x = 0.347

OpenStudy (anonymous):

my guess \[e^x-2e^{-x}=1\\ x=\ln(2)\]

OpenStudy (perl):

did you try to factor it ?

OpenStudy (anonymous):

i do not know how to do this at all

OpenStudy (perl):

\[e ^{x}-2e^{-x}=1\]\[e ^{x}-2e^{-x}-1=0\]

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

x = 0.347 is this it

OpenStudy (perl):

nope, but as sattellite suggested, if you plug in x = ln 2 , that works

OpenStudy (jhannybean):

\[e^x -2e^{-x} = 1\]\[\frac{e^x}{e^x} (e^x -2e^{-x})=1\]\[\frac{1}{e^x} (e^{2x} -2)=1\]This way might work as well, no?

OpenStudy (jhannybean):

basically I factored out an \(e^{-x}\) which can be rewritten as \(\dfrac{e^x}{e^x}\)

OpenStudy (anonymous):

o ok so it is B

OpenStudy (jhannybean):

Then I multiplied the "top" by \(e^x\)

OpenStudy (perl):

yes

OpenStudy (perl):

@Jhannybean did you get x = ln 2 as your solution ?

OpenStudy (jhannybean):

I haven't solved it yet :P

OpenStudy (perl):

oh :)

OpenStudy (jhannybean):

\[\frac{1}{e^x} (e^{2x} -2)=1\]\[\frac{e^{2x}}{e^x} -\frac{2}{e^x} = 1\]\[e^{2x}-2=e^x\]I'm not quite sure how to solve this

OpenStudy (perl):

ohh i see now

OpenStudy (perl):

i misspoke earlier, you might be able to factor this

OpenStudy (jhannybean):

Oh... \[e^{2x} -e^x =2\]\[e^x = 2\]\[\log_e (e^x) = \log_e (2)\]\[x= \log_e (2) = \ln (2)\]

OpenStudy (jhannybean):

@satellite73 is this how you would have solved this?

OpenStudy (anonymous):

A is it right

OpenStudy (jhannybean):

input \(\ln (2)\) into your calculator, what do you get?

OpenStudy (jhannybean):

Yeah that's the same thing I did :P

OpenStudy (perl):

ok good we got the same answer :)

OpenStudy (anonymous):

start with \[e^{2x}-2=e^x\] and then \[e^{2x}-e^x-2=0\] treat it as a quadratic

OpenStudy (anonymous):

so A is right

OpenStudy (anonymous):

no it is c

OpenStudy (perl):

e^x - 2e^(-x) = 1 e^x - 2/e^x = 1 multiply both sides by e^x, which will cancel the e^x in the denominator e^x*e^x - 2 = e^x e^(x+x) - e^x -2 = 0 e^(2x) - e^x - 2 = 0 (e^x)^2 - 1*(e^x) -2 = 0 (e^x - 2 )*( e^x + 1 ) = 0 e^x - 2 = 0 , e^x + 1 = 0 e^x = 2 , e^x = -1 (no solution) ln(e^x) = ln 2 x = ln 2